# Provable Privacy with Non-Private Pre-Processing

### v1

@@ C.3 Sensitivity analysis of PCA
Now we show ${\mathbb{P}}\left[{{\mathcal{A}}_{\mathrm{PTR}}(S_{1})\neq\perp}\right]\leq\nicefrac{{\delta}}{{2}}$.
@@
Lemma M (RDP to Approximate DP [Mironov, 2017]).
@@
Lemma N (Basic composition of differential privacy Dwork and Roth [2014]).
@@
To bound this term, we upper bound $\left|{\lambda_{k}(\tilde{\Sigma})-\lambda_{k}(\Sigma)}\right|$ for any $k$. Using Weyl’s inequality [Schindler, 2015]
@@
We have shown that Algorithm 1 does not output $\perp$ with probability $1-\frac{1}{2\delta}\exp\left({\frac{\beta-\delta_{k}(S)}{12.2}n\varepsilon}\right)$. It remains to derive the convergence guarantee of Algorithm 1 when the output is not $\perp$. To show this, we utilize the property that PCA transforms the dataset to have a low-rank covariance matrix, which allows us to apply the dimension-independent convergence guarantee of DP-GD following the analysis of Song et al. [2021] and establish the desired convergence guarantee.
@@
By Theorem 3.1 in Song et al. [2021], part $(b)$ is upper bounded by $O\left(\frac{L\sqrt{1+k\log\frac{1}{\delta}}}{\varepsilon n}\right)$.
Lemma O (Theorem 3.1 in Song et al. [2021]).

### v4

@@ C.4 Sensitivity Analysis of Scaling
Now we show ${\mathbb{P}}\left[{{\mathcal{A}}_{\mathrm{PTR}}(S_{1})\neq\perp}\right]\leq\nicefrac{{\delta}}{{2}}$. Given a input dataset $S$, we denote the $\Gamma$ in step 1 of Algorithm 1 by $\Gamma(S)$.
@@
Lemma M (RDP to Approximate DP (Mironov, 2017)).
@@
Lemma N (Basic composition of differential privacy Dwork & Roth (2014)).
@@
To bound this term, we upper bound $\left|{\lambda_{k}(\tilde{\Sigma})-\lambda_{k}(\Sigma)}\right|$ for any $k$. Using Weyl’s inequality (Schindler, 2015)
@@
We have shown that Algorithm 1 does not output $\perp$ with probability $1-\frac{1}{2\delta}\exp\left({\frac{\beta-\delta_{k}(S)}{12.2}n\varepsilon}\right)$. It remains to derive the convergence guarantee of Algorithm 1 when the output is not $\perp$. To show this, we utilize the property that PCA transforms the dataset to have a low-rank covariance matrix, which allows us to apply the dimension-independent convergence guarantee of DP-GD following the analysis of Song et al. (2021) and establish the desired convergence guarantee.
@@
By Theorem 3.1 in Song et al. (2021), part $(b)$ is upper bounded by $O\left(\frac{L\sqrt{1+k\log\frac{1}{\delta}}}{\varepsilon n}\right)$.
Lemma O (Theorem 3.1 in Song et al. (2021)).
@@
Appendix F Experiment setups and discussion
F.1 Experiment setups
Data generation
The synthetic data is generated with the make_classification function in the sklearn library. We generate a 2-class low rank dataset consisting of 1000 data points with dimension 6000 and rank 50. We set the parameter n_cluster_per_class in make_classification to 1.
Models and allocation of privacy budget
We compared the excess empirical loss of three models in Figure 3. We provide the details of the models below.
  • Pre-processed DP pipeline: We employ non-private PCA to reduce the dimensionality of the original dataset to $k$ and then apply private logistic regression. In particular, we use the make_private_with_epsilon method from the Opacus library with PyTorch SGD optimizer with learning rate 1e-2, max_grad_norm = 10 and epochs = 10. The privacy parameters epsilon and delta are obtained by adjusting the desired overall privacy level with our framework (Theorem 17).
  • No pre-processing: We directly apply private logistic regression with the same parameters on the original high-dimensional dataset.
  • DP-PCA: We first implement the DP-PCA in Chaudhuri et al. (2012) and then apply private logistic regression. We allocate half of the privacy budget to DP-PCA and the remaining half to private logistic regression.
F.2 Discussion on clipping
In practice, when the sensitivity of the original function is unbounded, one can apply the technique of clipping to restrict the sensitivity (Abadi et al., 2016; Liu et al., 2022). We can also incorporate clipping into our pre-processed DP pipeline. Specifically, we first non-privately compute the PCA projection matrix $A_{k}$ with the original dataset $S$. We clip the original data to $S_{\text{clipped}}$ with the clipping threshold $C\in[0.1R,0.7R,0.99R]$, where $R$ represents the maximum norm of the original dataset. Then, we apply the projection matrix to the clipped dataset to obtain the pre-processed dataset $S_{\text{preprocessed}}=A_{k}A_{k}^{\top}S_{\text{clipped}}$. Finally, we apply private logistic regression on $S_{\text{preprocessed}}$, with the privacy parameter set by our framework.
As shown in Table 4, the effect of clipping on accuracy depends on the clipping threshold. This is because clipping reduces the $L_{2}$ sensitivity of PCA from $\nicefrac{{12R}}{{n\delta_{\min}^{k}}}$ to $\nicefrac{{12C}}{{n\delta_{\min}^{k}}}$, allowing for selection of a larger privacy parameter during private learning and better accuracy. However, clipping also introduces inaccuracies in the dataset.
Table 4: Excess empirical loss of clipped pre-processed DP-pipeline
Clipping 0.1 Clipping 0.7 Clipping 0.99 Pre-processed DP pipeline
1.0 0.84 0.89 0.88 0.89
2.0 0.84 0.89 0.89 0.89
5.0 0.85 0.90 0.90 0.90
However, we note that $R$ is dataset dependent and might lead to additional privacy leakage.